# Post-measurement Quantum Monte Carlo

### v1

@@ VI Discussion
In this work we have developed a versatile new QMC technique for the study of measurement-induced quantum phenomena. Our focus was on the effects of measurements on thermal states of many-body systems. At low (or vanishing) temperatures, the strong correlations in these states can amplify the nonlocal effects of local measurements. We implemented our technique to probe the effects of $\mathrm{SU}(2)$-symmetric measurements on Heisenberg AFMs on the square lattice. For this model we explored the role of the Anderson tower of states, the creation of SPT order via measurement, and the possibility of creating quantum states whose correlations realize exotic surface universality classes.
@@
The post-measurement SSE allows for the efficient evaluation of expectation values in measured quantum states when there is no QMC sign problem. Since the problem of sampling the outcomes of measurements from a quantum state can be solved by calculating such expectation values, our algorithm also allows for efficient sampling. At a high level, the idea is to expand the measured density matrix as a sum over large numbers of ‘strings’ (products) of local operators, and to express expectation values in the measured density matrix as averages over contributing operator strings. For large classes of density matrices, this decomposition can be performed in such a way that the averages can be evaluated over the steps of a Markov chain, i.e. via Monte Carlo sampling of operator strings. The existence of a decomposition with this property amounts to the absence of a ‘sign problem’.
@@
The correlations which can be probed in experiment are nonlinear functions of post-measurement density matrices, e.g. averages of $\braket{S^{z}_{j}}^{\bm{s}}\braket{S^{z}_{k}}^{\bm{s}}$ and the entanglement entropy $-\text{Tr}[\rho^{\bm{s}}\log\rho^{\bm{s}}]$ over observed outcomes $\bm{s}$. Nonlinear averages of this kind can have rather exotic behavior, with stark differences between different post-measurement states as discussed throughout this work. Notably, using techniques from conformal field theory (CFT), Ref. [25] recently identified one-dimensional critical ground states where these nonlinear correlation functions exhibit multifractal scaling. Exploring the statistical properties of measurement-averaged correlation functions using our algorithms is promising direction for future work.
@@ Acknowledgements.
Appendix A Inhomogeneous imaginary time
@@
where $L^{M}_{d}$ is the number of type $d$ operators in the string $\mathcal{M}^{\bm{s}}_{\bm{d}}$ indicated on the left. Note also that the matrix element $\braket{\alpha}{\mathcal{M}^{\bm{s}}_{\bm{d}}\mathcal{H}_{\bm{a}}}{\alpha}$ remains a boolean variable, i.e. it is unity for allowed configuration and zero otherwise. The post-measurement partition functions for singlet and triplet outcomes differ in the sign factors accompanying type $d=2$ operators, as well as in the configurations which vanish when acted on by $d=1$ operators.
@@
As usual, the periodicity in imaginary time is satisfied by $\mathcal{H}_{\bm{a}}$ and $\mathcal{M}^{\bm{s}}_{\bm{d}}$ operator strings together, i.e. |$\ket{\alpha(0)}=\ket{\alpha(L-1)}=\ket{\alpha}$, and we illustrate an example configuration in Figure 15. The diagonal and off-diagonal updates used to sample operator strings contributing to this partition function have the same form as described in Sec. III.3.
Figure 15: Example SSE configuration within the inhomogeneous imaginary time scheme for the $2\times 6$ lattice shown in Figure 5. Here $\bm{s}=\bm{0}$ and the operator string consists of $\mathcal{H}_{\bm{a}}$ and $\mathcal{M}^{\bm{0}}_{\bm{d}}$. The propagated state |$\ket{\alpha(6)}$, indicated by purple dashed line, separates the $\mathcal{H}_{\bm{d}}$ and $\mathcal{M}^{\bm{0}}_{\bm{d}}$ operator strings.
Appendix B Post-measurement expectation values within inhomogeneous imaginary time scheme
Following weak measurements of the kind described in Appendix A, the expectation value for an observable $O$ is given by

### v2

@@ VII Discussion
In this work we have developed a versatile new QMC technique for the study of measurement-induced quantum phenomena. Our focus was on the effects of measurements on thermal states of many-body systems. At low (or vanishing) temperatures, the strong correlations in these states can amplify the nonlocal effects of local measurements. First, we implemented our technique to probe the effects of $\mathrm{SU}(2)$-symmetric measurements on Heisenberg AFMs on the square lattice. For this model we explored the role of the Anderson tower of states, the creation of SPT order via measurement, and the possibility of creating quantum states whose correlations realize exotic surface universality classes. Although these measurements typically generate a QMC sign problem when working a site-local basis, there is no sign problem when performing single spin measurements in a fixed basis. Using such a scheme, and studying averaged properties of the resulting ensemble of post-measurement states, we showed numerically that $\mathrm{O}(3)$ quantum critical correlations are fragile to measurements of the order parameter.
@@
The post-measurement SSE allows for the efficient evaluation of expectation values in measured quantum states when there is no QMC sign problem. At a high level, the idea is to expand the measured density matrix as a sum over large numbers of ‘strings’ (products) of local operators, and to express expectation values in the measured density matrix as averages over contributing operator strings. For large classes of density matrices, this decomposition can be performed in such a way that the averages can be evaluated over the steps of a Markov chain, i.e. via Monte Carlo sampling of operator strings. The existence of a decomposition with this property amounts to the absence of a ‘sign problem’.
@@
The correlations which can be probed in experiment are nonlinear functions of post-measurement density matrices, e.g. averages of $\braket{S^{z}_{j}}^{\bm{s}}\braket{S^{z}_{k}}^{\bm{s}}$ and the entanglement entropy $\text{Tr}[\rho^{\bm{s}}\log\rho^{\bm{s}}]$ over observed outcomes $\bm{s}$. Nonlinear averages of this kind can have rather exotic behavior, with stark differences between different post-measurement states as discussed throughout this work. Notably, using techniques from conformal field theory (CFT), Ref. [25] recently identified one-dimensional critical ground states where these nonlinear correlation functions exhibit multifractal scaling. Exploring the statistical properties of measurement-averaged correlation functions using our algorithms is promising direction for future work.
@@
The data and code used to generate the plots in this paper are publicly available at Zenodo [69].
@@ Acknowledgements.
Appendix A Loop updates
Here we discuss the loop updates used for the QMC calculations throughout this work. These updates are crucial for achieving ergodicity and also for maintaining computational efficiency when sampling operator string configurations composed of both diagonal and off-diagonal Hamiltonian terms.
Recall that, in the SSE, the full set of operator strings acting on a basis state must return that state to itself, i.e. there are periodic boundary conditions in imaginary time. Inserting off-diagonal operators into the string while respecting these periodic boundary conditions is, at least naively, straightforward using local updates, such as pairwise insertions and deletions of these operators. Unfortunately, such updates typically fail to sample the full configuration space. This is because local updates cannot alter global topological features of the operator sequence, such as loop winding numbers.
Loop updates address this challenges by constructing and flipping non-local clusters of spins or operator segments. These updates satisfy detailed balance while enabling transitions between topological sectors, thereby ensuring that all relevant operator string configurations are accessible. In practice, a full loop update sweep visits each operator, on average, once, resulting in a computational cost that scales linearly with both the number of lattice sites NN and the inverse temperature $\beta$, that is, $\mathcal{O}(\beta N)$.
In the parameter regimes explored in this work, we observe no signs of critical slowing down. Loop updates remain effective even in the presence of long-range correlations and near-critical behavior. To ensure convergence to equilibrium, we follow standard QMC procedures by monitoring observables as a function of $\beta$. We find that choosing $\beta\propto 3N$ is sufficient to obtain accurate results for all observables of interest, despite the vanishing finite-size gaps associated with the Anderson tower of states (discussed in Sec. IV).
Appendix B Inhomogeneous imaginary time
@@
where $L^{M}_{d}$ is the number of type $d$ operators in the string $\mathcal{M}^{\bm{s}}_{\bm{d}}$ indicated on the left. Note also that the matrix element $\braket{\alpha|\mathcal{M}^{\bm{s}}_{\bm{d}}\mathcal{H}_{\bm{a}}|\alpha}$ remains a boolean variable, i.e. it is unity for allowed configuration and zero otherwise. The post-measurement partition functions for singlet and triplet outcomes differ in the sign factors accompanying type $d=2$ operators, as well as in the configurations which vanish when acted on by $d=1$ operators.
@@
[…]
Appendix C Post-measurement expectation values within inhomogeneous imaginary time scheme
Following weak measurements of the kind described in Appendix B, the expectation value for an observable $O$ is given by